1 To 2 2. X = 1 1 -2 X 2 2 -1linear differential equations
Answer #1

oll 2 X = -2 X 2-1 slun we find eigen value of this matrik 2 -2 2 Biost C-a 1-0 ? -12 -a [-(i-a)(1+2) +4] +2 [+4+(1+a]41[a=24b) |-反了 2 TW 小。 -+反」 反义-2 २ -十五」 只 १ - 1+2」 45 b 13:00 = 一 -ELE (e){+反 「4-C? 「如 Lu 一 200 - y P I us|PC2 - V%z 反對大(24) 12 ***) (f) + Hat (3) ( C 2 D 2